# Overview of 3GPP Release 19 Study on Channel Modeling Enhancements to TR 38.901 for 6G

### v1

@@ I Introduction
In this paper, we present a summary of the key enhancements, along with an overview of the underlying rationale, and key technical discussions that guided the validation, refinement, and enhancements of [1], that are incorporated into [2]. Section II presents the scope of the 3GPP Rel-19 SI, followed by Section III, which discusses the various channel modeling components studied and incorporated [2]. These include the addition of a new SMa scenario (Subsection III-A), realistic UT antenna models (Subsection III-B), refinements and new additions to channel model parameters (Subsection III-C), a framework for modeling variability in power among different polarizations (Subsection III-D), NF propagation (Subsection III-E), and SNS (Subsection III-F). Finally, conclusions are presented in Section IV.
@@
  • Integration of SNS effect modeling to account for antenna element-wise power variation at the both the BS and UT.
TABLE I: Measurements (M) and Simulation (S) results provided in 3GPP Rel-19 SI for different environments (Env.), Frequencies (Freq.) and Bandwidths (BW) by various sources for validating, refining and enhancing the 3GPP TR 38.901 channel models.
@@
ZTE S 7, 8.4 NA NF impact on delay, power, AOA, AOD, ZOA, ZOD, phase and
SNS impact on power
BUPT, Spark M 13 0.4 DS, ASD, ASA, ZSA, ZSD, Number of Cluster, Number of Rays/Cluster, Cluster - DS, ASD, ASA, ZSA, shadowing;
Keysight M 10.1 0.5 PL, DS, Recien K Factor, XPR, ASA, ZSA, ASD, ZSD
Ericsson M 0.8625, 2.011, 5.02, 10.297, 22.001, 37 Conitunous Wave (CW) Excess PL
@@
M 10 0.5 DS, ASD, ASA, ZSD, number of cluster, Absolute time of arrival
@@
BUPT, Spark M 13 0.4 DS, ASD, ASA, ZSA, ZSD, Number of Cluster, Number of Rays/Cluster, Cluster - DS, ASD, ASA, ZSA, shadowing; NF and SNS
@@
M 1.8775, 5.25 CW, 0.2 XPR, fast fading characteristics
@@
M 2.6, 13 20, 0.0055 Absolute time of arrival, Polarisation, XPR
@@
Apple S 8 NA PL, DS, ASD, ASA, ZSA, Number of Cluster
@@
Rohde & Schwarz M 14 2 PL, DS, AS, Number of Multipaths, Ricean K-factor
@@
ZTE M 6-10 NA DS, Near field impact on delay
BUPT, Spark M 13 0.4 DS, ASD, ASA, ZSA, ZSD, Number of Cluster, Number of Rays/Cluster, Cluster - DS, ASD, ASA, ZSA; NF and SNS
Keysight M 10.1 0.5 PL, DS, Recian K Factor, XPR, ASA, ZSA, ASD, ZSD
@@
Huawei, HiSilicon M/S 10, 6.7 0.5, NA DS, ASD, ASA, ZSD, Number of Cluster, Absolute time of arrival, SNS
@@ III-B UT Antenna Modeling
A model for the BS antenna panels is provided in Section 7.3 [1]. In contrast, on the UT side, antenna modeling has historically remained oversimplified. [1] considers only isotropic UT antenna patterns with single or dual cross-polar components. For advanced scenarios, such as MIMO [9] and FR2 studies [10], within 3GPP, more detailed UT antenna configurations were introduced. However, these assumptions remain unrealistic for handheld UTs due to:
@@
To address these limitations, a more realistic UT antenna modeling framework was introduced in 3GPP Rel-19 SI. To obtain the frequency dependent combined UT radiation pattern, the physical dimensions of the UT was introduced. A reference UT size of 15 cm × 7 cm for handheld devices and 20 cm × 20 cm for CPE which imposes a fixed distance between antennas was adopted. Furthermore, eight candidate antenna locations for handheld UTs and nine for CPEs were identified, as shown in Fig. 2 and Fig 7.3-6 [2], respectively. Each UT antenna is oriented along the vector from the device center to its location and the direction of maximum gain of the antenna are aligned with these directions. The new antenna pattern for the UT is defined in Table 7.3-2 [2] as shown in Fig. 2 on the left.
However, specifying only the maximum gain direction is insufficient to fully define the antenna radiation pattern orientation. Polarization must also be considered. When a single antenna field pattern is considered per location, the UT antenna reference radiated field is vertically polarized with all gain concentrated in the theta component. This configuration is referred to as the polarization direction along the vertical $Z^{\prime\prime}$ axis. When the reference radiation pattern is translated to the particular location, the maximum gain and the polarization directions are aligned as shown in Fig. 2 on the right, where dashed lines indicate gain direction and arrows indicate polarization.
Finally, the polarization components are transformed into the global coordinate system (GCS) based on the orientation of the UT itself, using the coordinate transformation procedures defined in Section 7.1 of [1]. As a result, both polarized receive field patterns $F_{rx,u,\theta}$ and $F_{rx,u,\phi}$ of UT antenna $u$ are generally non-zero, resulting in elliptical polarization. These components are used to compute the channel coefficients of the fast fading model described in section 7.5 of [1].
Similar considerations apply when a candidate location supports two orthogonal polarization field patterns.
To further improve the modeling accuracy, two new optional components allow simulation of power imbalance across UT antennas. The first model introduces a randomized loss per antenna port. This does not model the entire radio frequency (RF) chain, but accounts for variation in antenna performance due to placement, shape, and implementation differences. Although by default no imbalance is modeled, during 3GPP discussions, random sampling values per antenna ranging from -2 dB to + 3 dB were proposed[11].
@@
The second enhancement introduces a model to simulate SNS on the UT side, accounting for power variations due to blockages from the user’s hand or head. [1] already defines two blockage models: Model A, which uses a stochastic approach including self-blocking, and Model B, which adopts a geometric method to capture human and vehicular blockage. Both models address far-field blockage effects, applying attenuation equally across all UT antennas. However, Cellular Telecommunications and Internet Association (CTIA)-based phantom model [12] simulations show that uniform attenuation across blocked angles deviates from real single-hand grip behavior.
New UT SNS model assumes that 90% of UTs are affected by blockage, with the remaining 10% operating in free space. Three grip scenarios are modeled: single-hand grip, dual-hand grip, and head-and-hand grip (see Fig. 3). Their occurrence probabilities are defined in Table 7.6.14.2-1 of [2]. Each grip scenario affects the UT antennas differently. Attenuation values per antenna location are specified in Table 7.6.14.2-2, and are based on simulations for the frequencies up to 8.4 GHz and on measurements for around 15 GHz. Incorporating element-wise power variations due to SNS enables a more accurate performance evaluation of massive MIMO systems.
@@ III-C Channel Model Parameters
  • The modeling of absolute time of arrival for NLOS was extended across all scenarios based on extensive ray tracing simulations to support sensing and positioning use cases.
Additionally, channel sparsity increases with frequency, resulting in fewer dominant multipath components and reduced channel rank. However, Table 7.5.6 [1] imposed fixed number of rays per cluster for narrow band systems and Eq. (7.6-8) [1] imposed a minimum of 20 rays per cluster for ELAAs, wideband system across all frequency bands and scenarios. This fixed and minimum lower bound did not accommodate the presence of fewer dominant multipath components as observed in measurements. To address this, Eq. (7.6-8) [2] was updated to allow modeling of a variable number of rays per cluster for ELAAs, wideband system, which can be less than 20 and frequency dependent for various scenarios.
TABLE III: Updated and Newly Introduced Channel (*) Model Parameters in 3GPP Rel-19 SI for various scenarios, applicable over the entire frequency range of 0.5–100 GHz. ”NA” indicates that the original values are retained.
@@ III-D Polarization
These assumptions implied a form of symmetry or balance between VP and HP, which influenced several radio interface design choices such as MIMO codebook construction. To validate these assumptions, a thorough re-examination of earlier measurements as well as new measurements of the polarization transformations between the transmit and receive antennas was conducted. While the earlier modeling assumptions were broadly accurate in terms of identical average powers among co and cross polarization components, with cross polarization components exhibiting weaker power compared to co polar components, the study revealed some variability around these assumed identical magnitudes in real-world propagation environments, which was not captured in the original model [1]. To account for the variability in power among the different polarizations, a polarization variability model was introduced in Section 7.6.16 [2].
@@ III-E Near Field Channel Model
The deployment of ELAAs is a key enabler for enhanced mobile broadband, improved spectral efficiency, and increased capacity in next-generation MIMO systems. As the physical aperture of the antenna arrays grows, the propagation transitions from far field to NF. Far field propagation assumes that the wavefront is planar whereas, NF propagation models the wavefront as spherical. The channel models in [1] are based on planar wavefront (far field) and therefore do not accurately capture the characteristics of spherical wavefront. This limitation significantly reduces the applicability of the channel models [1] for ELAAs deployments, where accurate modeling of NF propagation is essential. Thus, to address this limitation and accurately model NF propagation, several key aspects must be considered:
@@ III-F Modeling of Spatial Non-Stationarity
Figure 4: Near-field (NF) propagation and Spatial non-stationarity (SNS) at the BS and UT.
@@
For modeling SNS at the BS, two alternative approaches are proposed - a physical blocker-based model and a stochastic model. In the physical blocker-based model, the number, locations and dimensions of the blockers are explicitly determined as per Table 7.6.14.1.1-1 [2]. Based on the spatial relationship between clusters, blockers, and antenna elements, the antenna element-wise blockage conditions and power attenuation factors are determined based on Eq. (7.6-52) [2]. In contrast, the stochastic model uses a statistical approach. In this approach, every UE is assigned a visibility probabilty that governs the fraction of its clusters impacted by SNS. Next, for each impacted cluster a visibility region is defined that determines the fraction of the base station antennas that are visible to this cluster. The clusters or rays affected by SNS are determined using Eq. (7.6-53, 7.6-54) [2], their corresponding visible regions on the antenna array are computed using Eq. (7.6-57) [2], and the antenna element-wise power attenuation factors in Eq. (7.6-58) [2] are randomly generated based on statistical distributions derived data. Additionally, the modeling of SNS at UT is described earlier in Section III-B.
@@ IV Conclusion
This paper provides an overview of the 3GPP Rel 19 standardization efforts on channel modeling for 6G. These include the introduction of a new SMa scenario along with its associated modeling components, incorporation of realistic antenna models for handheld UTs, support for a variable number of clusters, reduced rays per cluster, and variability in power among different polarization. Additionally, a new plywood material penetration model, and a low loss outdoor to indoor penetration model for SMa scenario was introduced. Absolute delay modeling was extended to cover scenarios such as UMa, UMi, RMa, InH, and SMa. Moreover, channel parameters such as DS, ASA, ASD, and ZSA for UMi and UMa scenarios were updated based on new measurements and simulations. Also, refinements were made to cluster ASD for UMa LOS, NLOS, and O2I scenarios. Furthermore, channel modeling for NF propagation and SNS was developed for evaluating and analyzing future MIMO systems employing ELAA.

### v2

@@ I Introduction
In this paper, we present a summary of the key enhancements, along with an overview of the underlying rationale, and key technical discussions that guided the validation, refinement, and enhancements of [1], that are incorporated into [2]. Section II presents the scope of the 3GPP Rel-19 SI, followed by Section III, which discusses the various channel modeling components studied and incorporated [2]. These include the addition of a new SMa scenario (Subsection III-A), realistic UT antenna models (Subsection III-B), refinements and new additions to channel model parameters (Subsection III-C), a framework for modeling variability in power among different polarizations (Subsection III-D), NF propagation (Subsection III-E), and SNS effect (Subsection III-F). Finally, conclusions are presented in Section IV.
@@
  • Integration of SNS effect modeling to account for antenna element-wise power variation at both the BS and UT.
TABLE I: Measurements (M) and Simulation (S) results provided in 3GPP Rel-19 SI for different environments (Env.), Frequencies (Freq.) and Bandwidths (BW) by various sources for validating, refining and enhancing the 3GPP TR 38.901 channel models. “NA” indicates that the values were not reported. “CW” indicates continuous wave.
@@
ZTE S 7, 8.4 NA NF, SNS
BUPT, Spark M 13 0.4 DS, ASD, ASA, ZSA, ZSD, number of clusters, number of rays per cluster, cluster: DS, ASD, ASA, ZSA, shadowing
Keysight M 10.1 0.5 PL, DS, Rician K-factor, polarization, ASA, ZSA, ASD, ZSD
Ericsson M 0.8625, 2.011, 5.02, 10.297, 22.001, 37 CW Excess PL
@@
M 10 0.5 DS, ASD, ASA, ZSD, number of clusters, absolute time of arrival
@@
BUPT, Spark M 13 0.4 DS, ASD, ASA, ZSA, ZSD, number of clusters, number of rays per cluster, cluster: DS, ASD, ASA, ZSA, shadowing; NF, SNS
@@
M 1.8775, 5.25 CW, 0.2 Polarization, fast fading characteristics
@@
M 2.6, 13 20, 0.0055 Absolute time of arrival, polarization
@@
Apple S 8 NA PL, DS, ASD, ASA, ZSA, number of clusters
@@
Rohde & Schwarz M 14 2 PL, DS, AS, number of rays, Rician K-factor
@@
ZTE M 6-10 NA DS, NF
BUPT, Spark M 13 0.4 DS, ASD, ASA, ZSA, ZSD, number of clusters, number of rays per cluster, cluster: DS, ASD, ASA, ZSA; NF and SNS
Keysight M 10.1 0.5 PL, DS, Rician K-factor, polarization, ASA, ZSA, ASD, ZSD
@@
Huawei, HiSilicon M/S 10, 6.7 0.5, NA DS, ASD, ASA, ZSD, number of clusters, absolute time of arrival, SNS
@@ III-B UT Antenna Modeling
A model for the BS antenna panels is provided in Section 7.3 [1]. In contrast, on the UT side, antenna modeling has historically remained oversimplified. [1] considers only isotropic UT antenna patterns with single or dual cross-polar components. For advanced scenarios, such as multiple-input multiple-output (MIMO) [9] and FR2 studies [10], within 3GPP, more detailed UT antenna configurations were introduced. However, these assumptions remain unrealistic for handheld UTs due to:
@@
To address these limitations, a more realistic UT antenna modeling framework was introduced in 3GPP Rel-19 SI. To obtain the frequency dependent combined UT radiation pattern, the physical dimensions of the UT was introduced. A reference UT size of 15 cm × 7 cm for handheld devices and 20 cm × 20 cm for CPE which imposes a fixed distance between antennas was adopted. Furthermore, eight candidate antenna locations for handheld UTs and nine for CPEs were identified, as shown in Fig. 2 and Fig 7.3-6 [2], respectively. Each UT antenna is oriented along the vector from the device center to its location and the direction of maximum gain of the antenna are aligned with these directions. The new antenna pattern for the UT is defined in Table 7.3-2 [2] and shown in Fig. 2 on the left. However, specifying only the maximum gain direction is insufficient to fully define the antenna radiation pattern orientation. Polarization must also be considered. When a single antenna field pattern is considered per location, the UT antenna reference radiated field is vertically polarized with all gain concentrated in the theta component. This configuration is referred to as the polarization direction along the vertical $Z^{\prime\prime}$ axis. When the reference radiation pattern is translated to the particular location, the maximum gain and the polarization directions are aligned as shown in Fig. 2 on the right, where dashed lines indicate gain direction and arrows indicate polarization. Finally, the polarization components are transformed into the global coordinate system based on the orientation of the UT itself, using the coordinate transformation procedures defined in Section 7.1 [2]. As a result, both polarized receive field patterns $F_{rx,u,\theta}$ and $F_{rx,u,\phi}$ of UT antenna $u$ are generally non-zero, resulting in elliptical polarization. These components are used to compute the channel coefficients of the fast fading model described in section 7.5 [2]. Similar considerations apply when a candidate location supports two orthogonal polarization field patterns.
To further improve the modeling accuracy, two new optional components allow simulation of power imbalance across UT antennas. The first model introduces a randomized loss per antenna port. This does not model the entire radio frequency chain, but accounts for variation in antenna performance due to placement, shape, and implementation differences. Although by default no imbalance is modeled, during 3GPP discussions, random sampling values per antenna ranging from -2 dB to + 3 dB were proposed[11].
@@
The second enhancement introduces a model to simulate SNS effect at the UT, accounting for power variations due to blockages from the user’s hand or head. [1] already defines two blockage models: Model A, which uses a stochastic approach including self-blocking, and Model B, which adopts a geometric method to capture human and vehicular blockage. Both models address far-field blockage effects, applying attenuation equally across all UT antennas. However, Cellular Telecommunications and Internet Association (CTIA)-based phantom model [12] simulations show that uniform attenuation across blocked angles deviates from real single-hand grip behavior. New UT SNS model assumes that 90% of UTs are affected by blockage, with the remaining 10% operating in free space. Three grip scenarios are modeled: single-hand grip, dual-hand grip, and head-and-hand grip as shown in Fig. 3. Their occurrence probabilities are defined in Table 7.6.14.2-1 [2]. Each grip scenario affects the UT antennas differently. Attenuation values per antenna location are specified in Table 7.6.14.2-2 [2] based on simulations for frequencies up to 8.4 GHz and based on measurements for frequencies around 15 GHz. Incorporating antenna element-wise power variations due to SNS effect enables a more accurate performance evaluation of massive MIMO systems.
@@ III-C Channel Model Parameters
  • The modeling of absolute time of arrival for non-line of sight (NLOS) was extended across all scenarios based on extensive ray tracing simulations to support sensing and positioning use cases.
Additionally, channel sparsity increases with frequency, resulting in fewer dominant multipath components and reduced channel rank. However, Table 7.5.6 [1] imposed fixed number of rays per cluster for narrow band systems and Eq. (7.6-8) [1] imposed a minimum of 20 rays per cluster for ELAAs and/or wideband system across all frequency bands and scenarios. This fixed and minimum lower bound for the number of intra cluster rays did not accommodate the presence of fewer dominant multipath components as observed in measurements. To address this, Eq. (7.6-8) [2] was updated to allow modeling of a variable number of rays per cluster for ELAAs and/or wideband system, which can be less than 20 and frequency dependent for various scenarios.
TABLE III: Updated and Newly Introduced (*) Channel Model Parameters in 3GPP Rel-19 SI for various scenarios, applicable over the entire frequency range of 0.5–100 GHz. “NA” indicates that the original values are retained.
@@ III-D Polarization
These assumptions implied a form of symmetry or balance between VP and HP, which influenced several radio interface design choices such as MIMO codebook construction. To validate these assumptions, a thorough re-examination of earlier measurements as well as new measurements of the polarization transformations between the transmit and receive antennas was conducted. While the earlier modeling assumptions were broadly accurate in terms of identical average powers among co and cross polarization components, with cross polarization components exhibiting weaker power compared to co polar components, the study revealed some variability around these assumed identical magnitudes in real-world propagation environments, which was not captured in the original model in Eq. (7.5-28) [1]. To account for the variability in power among the different polarizations, a polarization variability model was introduced in Section 7.6.16 [2].
@@ III-E Near Field Channel Model
The deployment of ELAAs is a key enabler for enhanced mobile broadband, improved spectral efficiency, and increased capacity in next-generation MIMO systems. As the physical aperture of the antenna arrays grows, the propagation transitions from far field to NF. Far field propagation assumes that the wavefront is planar whereas, NF propagation models the wavefront as spherical. The channel models in [1] are based on planar wavefront (far field) and therefore do not accurately capture the characteristics of spherical wavefront (NF). This limitation significantly reduces the applicability of the channel models in [1] for ELAAs deployments, where accurate modeling of NF propagation is essential. Thus, to address this limitation and accurately model NF propagation, several key aspects must be considered:
@@ III-F Modeling of Spatial Non-Stationarity
Figure 4: NF propagation and SNS effect at both the BS and UT.
@@
For modeling SNS effect at the BS, two alternative approaches are proposed - a physical blocker-based model and a stochastic model. In the physical blocker-based model, the number, locations and dimensions of the blockers are explicitly determined as per Table 7.6.14.1.1-1 [2]. Based on the spatial relationship between clusters, blockers, and antenna elements, the antenna element-wise blockage conditions and power attenuation factors are determined based on Eq. (7.6-52) [2]. In contrast, the stochastic model uses a statistical approach. In this approach, every UE is assigned a visibility probability that governs the fraction of its clusters impacted by SNS. Next, for each impacted cluster a visibility region is defined that determines the fraction of the BS antennas that are visible to this cluster. The clusters or rays affected by SNS are determined using Eq. (7.6-53, 7.6-54) [2], their corresponding visible regions on the antenna array are computed using Eq. (7.6-57) [2], and the antenna element-wise power attenuation factors in Eq. (7.6-58) [2] are randomly generated based on statistical distributions derived from measurement and simulation data. Additionally, the modeling of SNS effect at the UT is described earlier in Section III-B.
@@ IV Conclusion
This paper provides an overview of the 3GPP Rel-19 standardization efforts on channel modeling for 6G. These include the introduction of a new SMa scenario along with its associated modeling components, incorporation of realistic antenna models for handheld UTs, support for modeling variable number of clusters, reduced rays per cluster, and variability in power among different polarizations. Additionally, a new plywood material penetration loss model, and a low loss O2I building penetration loss model for SMa scenario was introduced. Absolute delay modeling was extended to cover scenarios such as UMa, UMi, RMa, InH, and SMa. Moreover, channel parameters such as DS, ASA, ASD, and ZSA for UMi and UMa scenarios were updated based on new measurements and simulations. Also, refinements were made to cluster ASD for UMa LOS, NLOS, and O2I scenarios. Furthermore, channel modeling for NF propagation and SNS effect was developed for evaluating and analyzing future MIMO systems employing ELAA.